OpenStudy (anonymous):

ABCD is a rectangle. If AC=3y/5 and BD=3(y-4). find the length of each diagonal

OpenStudy (anonymous):



OpenStudy (anonymous):

Now, diagonals of a rectangle are equal so AC=BD

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

AC=BD 3y/5 =3(y-4) y/5=y-4 y=5y-20 y=5

OpenStudy (anonymous):

Now, can u?